OpenStudy (anonymous):

rational expression to lowest terms ... help please, 6m^2/9m^4-15m^3

OpenStudy (anonymous):

\[\frac{6m^2}{9m^{4} - 15m^{3}}\] They have a common factor:\[\frac{3m^{2} * 2}{3m^{2} *(3m^{2} - 5m^{1})}\] Cancel out the \[3m^{2}\] so you get \[\frac{2}{(3m^{2} - 5m^{1})}\] I hope that's what you're looking for. If not, sorry!

OpenStudy (anonymous):

thats perfect I had the numerator reversed on my paper I had (m+3)(m+2) so I did the numerator wrong ! thank you soo much I was re writting this for 30 min now haha